Box of mass 20 kilograms is kept on a smooth horizontal surface. Object of mass 16 kilograms is kept on the box. The two objects are connected by a light in extensible string and the string goes over smooth Pulley. A force of 6 Newton is applied on the box, the acceleration of the box?

1
Expert's answer
2022-05-31T16:11:21-0400

"a=\frac F{m_2-m_1}=1.5~\frac N{kg}."